A bag contains 10 balls, out of which 2 are white, 5 are red and 3 are black balls. If three are drawn at random, then the probability that two balls are of same colour is

 

Grade:12

2 Answers

Arun
4 years ago
 
Dear student
 
ways of drawing three balls out of which 2 balls are same = 
 = 2c2 *8c1 + 5c2 * 5c1 + 3c2 * 7c1 = 8 + 50 + 21 = 79
 
hence probability = 79/10c3 = 79/120
 
Hope it helps
Vikas TU
4 years ago
Dear student 
Please follow this link for generalised solutin of such types of problems 
Hope this will help 
Good Luck